OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation.

OpenStudy (hoblos):

let's find the slope first the slope between two points (x1,y1) ans (x2,y2) is \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] so what's the slope between (2, 3) and (6, 8)

OpenStudy (anonymous):

5/4?

OpenStudy (hoblos):

correct now we have to apply this formula to get the equation \[y - y_1 = m(x - x_1)\]

OpenStudy (anonymous):

Is it y - 3 =5/4 (x-2) ?

OpenStudy (hoblos):

correct

OpenStudy (anonymous):

thank you!

OpenStudy (hoblos):

any time :)